A couple received a $128,000 inheritance the year they turned 48 and invested it in a fund that earns 6.4% compounded semiannually. If this amount is deferred for 14 years (until they retire), how much will it provide at the end of each half year (in dollars) for the next 20 years after they retire? (Round your answer to two decimal places.)
Answer #1
Semi annual cash flow = Present value / Present value of annuity of 1
= $ 1,28,000.00 / 18.31345
= $       6,989.40
Working:
Present value of annuity of 1 = (1-(1+i)^-n)/i Where,
= (1-(1+0.032)^-28)/0.032 i = 6.4%/2 = 0.032
= 18.31344922 n = 14*2 = 28